OpenStudy (masterchief):

Create slope-intercept form with points that pass through (4,-7) and (-2,3)

OpenStudy (anonymous):

Find the slope first, which would be 10/-6 or -5/3. Then what you want to do is plug it into slope intercept form. This would be: y+7=-5/3(x-4). Does that sound right?

OpenStudy (masterchief):

Don't you just substitute the x and y? Shouldn't it be 7=-5/3(-4)

OpenStudy (anonymous):

After it would be reduced into slope intercept form, this is what I got: y=-5/3x -1/3

OpenStudy (masterchief):

I got a positive slope and positive b when i plugged in -2,3

OpenStudy (masterchief):

Well?

OpenStudy (anonymous):

Sorry, one second

OpenStudy (anonymous):

Okay, how did you find the slope?

OpenStudy (masterchief):

3-(-7) over -2-4

OpenStudy (anonymous):

slope=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

That would equal 10/-6, correct?

OpenStudy (masterchief):

Yes

OpenStudy (anonymous):

which reduces to -5/3

OpenStudy (masterchief):

Yup

OpenStudy (anonymous):

Pick either of the two points to use our forumla to get it down to slope intercept form. The formula we will use is: y-y1=m(x-x1)

OpenStudy (masterchief):

Oh what am i saying, i meant i got a positive b

OpenStudy (anonymous):

Oh, okay, do you understand the process though?

OpenStudy (masterchief):

Process of getting the slope or b?

OpenStudy (anonymous):

for getting the entire equation. Let me show you my reasoning. Lets pick the points (4,-7). When we plug them into our equation we get y+7=-5/3(x-4). We need to reduce this now, so let us distribute the -5/3. This will get us our new equation of y+7=-5/3x+20/3. Correct so far? Now we need to get rid of the 7 on that side so lets subtract. In order to do that we need common denominatiors. So we will end up have 21-20/3. The final answer is y=-5/3x+1/3

OpenStudy (anonymous):

Got my numbers mixed up before, but there it is =)

OpenStudy (masterchief):

That's what I got, what I did was I picked (4,-7), substituted them for x and y, solved the equation to get b, and plugged it in to y=-5/6x+b to get y=-5/6x+1/3